# Towards Sample-specific Backdoor Attack with Clean Labels via Attribute Trigger

### v2

@@ Abstract
Currently, sample-specific backdoor attacks (SSBAs) are the most advanced and malicious methods since they can easily circumvent most of the current backdoor defenses. In this paper, we reveal that SSBAs are not sufficiently stealthy due to their poisoned-label nature, where users can discover anomalies if they check the image-label relationship. In particular, we demonstrate that it is ineffective to directly generalize existing SSBAs to their clean-label variants by poisoning samples solely from the target class. We reveal that it is primarily due to two reasons, including (1) the ‘antagonistic effects’ of ground-truth features and (2) the learning difficulty of sample-specific features. Accordingly, trigger-related features of existing SSBAs cannot be effectively learned under the clean-label setting due to their mild trigger intensity required for ensuring stealthiness. We argue that the intensity constraint of existing SSBAs is mostly because their trigger patterns are ‘content-irrelevant’ and therefore act as ‘noises’ for both humans and DNNs. Motivated by this understanding, we propose to exploit content-relevant features, (human-relied) attributes, as the trigger patterns to design clean-label SSBAs. This new attack paradigm is dubbed backdoor attack with attribute trigger (BAAT). Extensive experiments are conducted on benchmark datasets, which verify the effectiveness of our BAAT and its resistance to existing defenses.
@@
The answer to the aforementioned question is in the negative. We argue that the intensity constraint of existing SSBAs is mostly because their trigger patterns are ‘content-irrelevant’ and therefore act as ‘noises’ for both humans and DNNs. Motivated by this understanding, in this paper, we propose to exploit content-relevant features, (human-relied) attributes, as the trigger patterns to design clean-label SSBAs. This new attack paradigm is dubbed backdoor attack with attribute trigger (BAAT). In general, our method is inspired by the decision process of humans. For example, we can use an adversary-defined hairstyle as our attribute trigger in facial recognition tasks. Since attribute is a high-level and complicated feature, the modifications between poisoned images (, the modified images containing trigger patterns) and their benign ones are sample-specific and can be large (, high intensity) while still preserving stealthiness. Their selection and design is also a feasible way to incorporate domain knowledge of the target task.
@@ 2.1 Backdoor Attacks
Backdoor attack is an emerging yet severe threat, revealing the training-phase security concerns of DNNs [8]. Specifically, the backdoored models behave normally on benign samples whereas their predictions will be maliciously changed whenever the adversary-specified trigger patterns appear. In this paper, we focus on poison-only backdoor attacks (, the adversaries can only modify the training dataset) in image classification. The backdoor threats with other threat models [16, 22, 23] or in other tasks [24, 25, 26, 27, 28] are out of our scope in this paper.
@@ 2.2 Backdoor Defenses
Pre-processing-based Defenses. These approaches pre-processed test images before feeding them into the model for prediction, motivated by the observations that backdoor attacks may lose effectiveness when the trigger used for attacking is different from the one used for poisoning [34, 13, 47]. These defenses are usually efficient since they did not require modifying the suspicious models.
Sample-filtering-based Defenses. These methods aim at filtering out poisoned samples. For example, defenders can identify malicious training samples based on their distinctive behaviors in the hidden feature space [48, 49, 50]. Recently, [51] proposed to filter poisoned testing samples by superimposing different images on the suspicious sample and observing their predictions. The smaller the prediction randomness, the more likely it is attacked. Most recently, Guo et al. [52] detected poisoned samples by analyzing their prediction consistency during pixel-wise amplification. The more consistent a sample, the more likely it is poisoned.
@@ 3.1 Design Clean-label Sample-specific Attacks by Poisoning Samples only from the Target Class
Settings. We conduct experiments on (a subset of) ImageNet dataset [53] having 100 random classes. Each class contains 500 images for training and 50 images for testing. We generalize the clean-label variants of WaNet and ISSBA (dubbed ‘WaNet-C’ and ‘ISSBA-C’, respectively) by poisoning samples only from the target class. Specifically, we set target class (, ‘n01443537’) and poison 80% samples from the target class. We conduct all attacks with both VGG-16 [54] and ResNet-18 [55], and implement them based on codes in BackdoorBox [56]. We use the default settings of ISSBA and adopt the settings of WaNet (without noise mode) where the kernel size is set as 32.
@@ 3.2.2 Sample-specific Triggers are More Difficult than Sample-agnostic Ones to Learn by DNNs
Settings. We compare ISSBA and WaNet with their sample-agnostic versions on the ImageNet subset with ResNet-18 under different poisoning rates. We randomly select three different poisoned samples generated by the standard ISSBA and exploit their pixel-wise differences to their benign version as trigger patterns to design three sample-agnostic versions of ISSBA (dubbed ’ISSBA-A (a)’, ’ISSBA-A (b)’, and ’ISSBA-A (c)’), respectively. We also design three sample-agnostic WaNets following the same setting.
@@
Theorem 1.
Settings. In this part, we conduct experiments on WaNet-C and ISSBA-C with different trigger intensities. Specifically, we set the intensity-related parameter of WaNet-C as and we amplify trigger perturbations of ISSBA-C with a factor from 0 to 8 (, ). Other settings are the same as those used in Section 3.1.
@@ 4.2 Backdoor Attack with Attribute Trigger (BAAT)
In general, attributes are the high-level features exploited by humans to describe and make predictions. However, it is difficult to provide a formal definition of the attribute, since the mechanism of the human visual system and the concept of features are very complicated and remain unclear. Luckily, we can at least find some suitable attributes in image classification tasks, based on some recent studies [59, 29, 60]. Here we used two representative tasks, , facial image and natural image recognition, as examples to describe how to design our attack with attribute triggers.
@@ 5.4.3 The Resistance to Trigger-synthesis-based Defenses
5.4.4 The Resistance to Pre-processing-based Defenses
Settings. We adopt a pre-trained auto-encoder trained on the ImageNet dataset for Auto-Encoder. Specifically, we first resize the images from to before feeding into the auto-encoder. After that, we shrink the pre-processed images back to , based on which to calculate the benign accuracy and the attack success rate; We implement ShrinkPad based on BackdoorBox, where the shrinking size is set to 12 pixels on both datasets.
5.4.5 The Resistance to Sample-filtering-based Defenses
@@ Acknowledgments
This work was partly done when Yiming Li was a research intern at Ant Group. We also sincerely thank Mr. Chengxiao Luo from Tsinghua University for his implementation of some preliminary experiments on the VGGFace2 dataset, and Prof. Yong Jiang from Tsinghua University and Dr. Haiqin Weng from Ant group for their valuable comments and suggestions on an early draft of this paper.
@@ Appendix A The Proof of Theorem 1
We have for the target poisoned sample since trigger patterns are additive. As such, for sample-specific samples: we have:, while for the sample-static samples: , where represents the backdoor trigger.
We treat our model as a k-way kernel least square classifier and use a cross-entropy loss for training the kernel. The output of is a k-dimensional vector. Let us assume be expected predictive confidences corresponding to the target class . Following previous works [58, 43], we know the kernel regression solution is:
@@
Accordingly, for sample-specific attacks, we have:
Similarly, for sample-agnostic attacks with the same and configurations, we have:
@@
where .
In particular, we know that since belongs to .
@@
![[Uncaptioned image]](extracted/5286442/YimingLi.png)
![[Uncaptioned image]](extracted/5286442/MingyanZhu.jpg)
![[Uncaptioned image]](extracted/5286442/junfeng.jpg)
![[Uncaptioned image]](extracted/5286442/TaoWei.jpg)
![[Uncaptioned image]](extracted/5286442/ShutaoXia.png)
![[Uncaptioned image]](extracted/5286442/ZhanQin.jpeg)

### v3

@@ Abstract
Currently, sample-specific backdoor attacks (SSBAs) are the most advanced and malicious methods since they can easily circumvent most of the current backdoor defenses. In this paper, we reveal that SSBAs are not sufficiently stealthy due to their poisoned-label nature, where users can discover anomalies if they check the image-label relationship. In particular, we demonstrate that it is ineffective to directly generalize existing SSBAs to their clean-label variants by poisoning samples solely from the target class. We reveal that it is primarily due to two reasons, including (1) the ‘antagonistic effects’ of ground-truth features and (2) the learning difficulty of sample-specific features. Accordingly, trigger-related features of existing SSBAs cannot be effectively learned under the clean-label setting due to their mild trigger intensity required for ensuring stealthiness. We argue that the intensity constraint of existing SSBAs is mostly because their trigger patterns are ‘content-irrelevant’ and therefore act as ‘noises’ for both humans and DNNs. Motivated by this understanding, we propose to exploit content-relevant features, (human-relied) attributes, as the trigger patterns to design clean-label SSBAs. This new attack paradigm is dubbed backdoor attack with attribute trigger (BAAT). Extensive experiments are conducted on benchmark datasets, which verify the effectiveness of our BAAT and its resistance to existing defenses. Our codes for reproducing main experiments are available at BackdoorBox and GitHub repository.
@@
The answer to the aforementioned question is in the negative. We argue that the intensity constraint of existing SSBAs is mostly because their trigger patterns are ‘content-irrelevant’ and therefore act as ‘noises’ for both humans and DNNs. Motivated by this understanding, in this paper, we propose to exploit content-relevant features, (human-relied) attributes, as the trigger patterns to design clean-label SSBAs. This new attack paradigm is dubbed backdoor attack with attribute trigger (BAAT). In general, our method is inspired by the decision process of humans. For example, we can use an adversary-defined hairstyle as our attribute trigger in facial recognition tasks. Specifically, the adversaries will first exploit a pre-trained attribute editor to assign the adversary-specified attribute of selected images as a particular value (without modifying their labels). These modified poisoned samples and remaining benign ones will be released to victims to train their models. Consequently, a model trained on these samples would misclassify any testing input, as long as its attribute value is changed to the adversary-specified one. Since attribute is a high-level and complicated feature, the modifications between poisoned images (, the modified images containing trigger patterns) and their benign ones are sample-specific and can be large (, high intensity) while still preserving stealthiness. Their selection and design is also a feasible way to incorporate domain knowledge of the target task.
@@ 2.1 Backdoor Attacks
Backdoor attack is an emerging yet severe threat, revealing the training-phase security concerns of DNNs [8]. Specifically, the backdoored models behave normally on benign samples whereas their predictions will be maliciously changed whenever the adversary-specified trigger patterns appear. In this paper, we focus on poison-only backdoor attacks (, the adversaries can only modify the training dataset) in image classification. The backdoor threats with other threat models [16, 22, 23] or in other tasks [24, 25, 26, 27, 28, 29, 30] are out of our scope in this paper.
@@ 2.2 Backdoor Defenses
Pre-processing-based Defenses. These approaches pre-processed test images before feeding them into the model for prediction, motivated by the observations that backdoor attacks may lose effectiveness when the trigger used for attacking is different from the one used for poisoning [36, 13, 49]. These defenses are usually efficient since they did not require modifying the suspicious models. Most recently, Xu et al. [50] proposed backdoor trigger inversion method that decouples benign instead of backdoor features to design a simple yet effective pre-processing-based defense.
Sample-filtering-based Defenses. These methods aim at filtering out poisoned samples. For example, defenders can identify malicious training samples based on their distinctive behaviors in the hidden feature space [51, 52, 53]. Recently, [54] proposed to filter poisoned testing samples by superimposing different images on the suspicious sample and observing their predictions. The smaller the prediction randomness, the more likely it is attacked. Most recently, [55, 56] detected poisoned samples by analyzing their input-level and weight-level prediction consistency. The more consistent a sample, the more likely it is poisoned.
@@ 3.1 Design Clean-label Sample-specific Attacks by Poisoning Samples only from the Target Class
Settings. We conduct experiments on (a subset of) ImageNet dataset [57] containing 100 random classes. Each class contains 500 images for training and 50 images for testing. We generalize the clean-label variants of WaNet and ISSBA (dubbed ‘WaNet-C’ and ‘ISSBA-C’, respectively) by poisoning samples only from the target class. Specifically, we set target class (, ‘n01443537’) and poison 80% samples from the target class. We conduct all attacks with both VGG-16 [58] and ResNet-18 [59], and implement them based on codes in BackdoorBox [60]. We use the default settings of ISSBA and adopt the settings of WaNet (without noise mode) where the kernel size is set as 32. We train the models with 30 epochs using a batch size of 128 and a learning rate of 0.001. The SGD optimizer is utilized with a momentum of 0.9 and a weight decay of .
@@ 3.2 Why Are Clean-label Sample-specific Backdoor Attacks Difficult to Succeed?
Unless otherwise specified, all settings are the same as those described in Section 3.1.
@@ 3.2.2 Sample-specific Triggers are More Difficult than Sample-agnostic Ones to Learn by DNNs
Settings. We compare ISSBA and WaNet with their sample-agnostic versions on the ImageNet subset with ResNet-18 under different poisoning rates. We randomly select three different poisoned samples generated by the standard ISSBA and exploit their pixel-wise differences to their benign version as trigger patterns to design three sample-agnostic versions of ISSBA (dubbed ‘ISSBA-A (a)’, ‘ISSBA-A (b)’, and ‘ISSBA-A (c)’), respectively. We also design three sample-agnostic WaNets following the same setting.
@@
Theorem 1.
Settings. In this part, we conduct experiments on WaNet-C and ISSBA-C with different trigger intensities. Specifically, we set the intensity-related parameter of WaNet-C as and we amplify trigger perturbations of ISSBA-C with a factor from 0 to 8 (, ).
@@ 4.2 Backdoor Attack with Attribute Trigger (BAAT)
In general, attributes are the high-level features exploited by humans to describe and make predictions. Arguably, attribute trigger is more effective mostly because it allows modifying images with a larger size and a higher intensity to dominate ground-truth features while still maintaining stealthiness. However, it is difficult to provide a formal definition of the attribute, since the mechanism of the human visual system and the concept of features are very complicated and remain unclear. Luckily, we can at least find some suitable attributes in image classification tasks, based on some recent studies [63, 31, 64]. In general, we can design effective attribute triggers via selecting unique attributes and modifying them with rare values but not abnormal. This ensures both the effectiveness and the stealthiness of our attack. Here we used two representative tasks, , facial image and natural image recognition, as examples to describe how to design our attack with attribute triggers.
@@ 5.4.3 The Resistance to Trigger-synthesis-based Defenses
5.4.4 The Resistance to Pre-processing-based Defenses
Settings. We adopt a pre-trained auto-encoder trained on the ImageNet dataset for Auto-Encoder. Specifically, we first resize the images from to before feeding into the auto-encoder. After that, we shrink the pre-processed images back to , based on which to calculate the benign accuracy and the attack success rate; We implement ShrinkPad based on BackdoorBox [60], where the shrinking size is set to 12 pixels on both datasets.
5.4.5 The Resistance to Sample-filtering-based Defenses
@@ 6 Discussions
6.1 The Analysis of Computational Complexity
In general, BAAT introduces only a small overhead during the one-time trigger generation phase, but it maintains comparable training and inference efficiency to clean models. We analyze the efficiency of our method as follows.
Attack Phase. In this phase, we employ an attribute editor to modify attributes and generate poisoned samples. Let and represent the size of the benign dataset and the poisoning rate, respectively. The computational complexity of our method is , as BAAT only requires poisoning a small number of randomly selected samples during this phase. While this introduces additional computation, it is a one-time operation performed prior to training. For instance, on the ImageNet dataset, our method generates poisoned samples for images at an average speed of 56.7 ms per image on an NVIDIA GeForce RTX 3080 Ti.
Training Phase. Incorporating poisoned samples either not change the training process nor increase the number of training samples. As such, the computational cost of this process remains identical to that of the standard one.
Inference Phase. During the inference process, BAAT does not require additional operations (, data augmentation or post-processing) of the attacked model. As a result, the inference time is the same as that of a clean model. Note that the adversary needs to generate the attacked image locally through the same pre-trained attribute editor used for poisoning, but its cost is negligible (, ).
@@ Acknowledgments
This research is supported in part by the National Key Research and Development Program of China under Grant 2021YFB3100300 and the National Natural Science Foundation of China under Grants (62171248, 62441238, 62072395, and U20A20178). This work was also partly done when Yiming Li was a research intern at Ant Group. We also sincerely thank Mr. Chengxiao Luo from Tsinghua University for his implementation of some preliminary experiments on the VGGFace2 dataset, and Prof. Yong Jiang from Tsinghua University and Dr. Haiqin Weng from Ant group for their valuable suggestions on an early draft of this paper.
@@ Appendix A The Proof of Theorem 1
We have for poisoned samples since trigger patterns are additive. As such, for sample-specific attacks, we have , while for the sample-agnostic attacks: , where represents the backdoor trigger.
We treat our model as a k-way kernel least square classifier and use a cross-entropy loss for training the kernel, and the output of is a k-dimensional vector. Let us assume be expected predictive confidences corresponding to the target class . Following previous works [62, 45], we know the kernel regression solution is:
@@
Accordingly, for sample-specific attacks and sample-agnostic attacks, we respectively have:
@@
where . In particular, we know that since belongs to .